• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name Jin Huang & Vince

    Log entry time 18:31:23 on October30,2009

    Entry number 296959

    This entry is a followup to: 296958

    keyword=Cosmic Check on LHRS

    Vince and I just had a quick check on the running cosmic run of 25138.

    As shown in the first plot, the HV of s1 PMTs are not well set. The problem is worse on PMT Left0 and Right5, whose signal is hardly recognizable from pedstal. There is no MIP peak on s2m ADC.

    TDC signal is generally OK as shown in 2nd plot. However, it's suspected that there are double pulsing in several channels of s1 TD. It's most noticeable in chan# Right0 and Left5.

    Vince plan to further investigate after trigger setup is stable and a pedstal run is taken.

    FIGURE 1

    FIGURE 2